# A Circuit-QED Lattice System with Flexible Connectivity and Gapped Flat Bands for Photon-Mediated Spin Models

### v1

@@ III.1 Transmon Qubits and Readout
III.1.1 Transmons in the Dispersive Limit
The transmon is the most common superconducting qubit currently in use [56, 9]. It consists of an anharmonic oscillator formed by a capacitor and a Josephson junction in parallel. The non-linearity of the Josephson junction [57] introduces a negative anharmonicity, which results in the |$\ket{0}\rightarrow\ket{1}$ transition frequency of the transmon, $\omega_{q}$, being higher in energy than all other transitions, allowing the transmon to be treated as an effective two-level system. If the single Josephson junction is replaced with a DC SQUID [57], the transition frequency of the transmon becomes tunable in-situ via applied magnetic flux. Unlike atomic systems, which can typically only be tuned a few tens of MHz with external magnetic fields, the transition frequency of a transmon can be varied over many GHz. The transmons featured in this work have a tuning range of approximately $2.5$ GHz - $10$ GHz, allowing them to be tuned into proximity with all of the modes in the photonic crystal.
Because of the low transition frequencies of transmon qubits, they cannot be read out via fluorescence or photon absorption, as would be standard for an atomic qubit. Instead, standard practice is to exploit off-resonant coupling to a microwave resonator with frequency $\omega_{r}$ to perform readout [9]. In the dispersive regime, where $|\omega_{q}-\omega_{r}|\gg g$, and approximating the transmon as a two-level system, the effective Hamiltonian takes the form:
@@ III.1.1 Transmons in the Dispersive Limit
where $\Delta=\omega_{q}-\omega_{r}$ and $\omega_{q}^{\prime}=\omega_{q}+g^{2}/\Delta$ is the Lamb-shifted qubit frequency. Crucially, the term $g^{2}\sigma_{z}a^{\dagger}a/\Delta$ constitutes a qubit-state dependent frequency shift of the resonator, known as the dispersive shift. Due to the native separation of energy scales between the coherent coupling strength $g$ of a transmon and the native decay rates, transmons can easily reach the so-called strong dispersive regime in which the dispersive shift is comparable to or exceeds the resonator linewidth $\kappa$ In this regime, the amplitude and phase of a monitor tone impinging on the resonator are strongly affected by the state of the qubit, and, as a result, homodyne or heterodyne measurements of this microwave tone yield QND measurements of the qubit state [9].
@@
Many geometries of resonator have been used in conjunction with transmons; however, in this work we concentrate only on the case of capacitive coupling to CPW resonators [45], because these resonators have the demonstrated capability to implement a large variety of unconventional band structures [18, 12]. These systems routinely display $g$’s in the range of $2\pi\times 50-2\pi\times 300$ MHz for single CPW resonators. In the case of an array of CPW resonators, such as the photonic crystal studied in this paper, the natural photonic degrees of freedom are normal modes, rather than individual cavity modes. In an array of $N$ resonators, a typical normal mode will be delocalized over the entire device, and due to the $N$-fold increase in the mode volume over that of a single isolated resonator, the light-matter coupling will become $g_{\mathrm{norm}}\sim g_{0}/\sqrt{N}$, where $g_{0}$ is the coupling between the qubit and the particular resonator in which it is located. Therefore, a CPW array with $\mathcal{O}(50-100)$ resonators can still have dispersive shifts which are an appreciable fraction of a MHz at a detuning of a GHz. Hence, the dispersive readout techniques which are normally used on transmon qubits in single-mode resonators can also be applied in resonator-array photonic crystal systems, using the normal modes.
III.1.2 Two-Tone Spectroscopy
The dispersive readout mechanism described above can be used as a spectroscopy technique for locating qubit transitions by combining two drive tones: one to excite the qubit and one to monitor the qubit state. In the case of conventional single-mode devices, the fixed-frequency monitor tone is applied on-resonance with a resonator, or a normal mode in the case of the CPW lattice device studied here, whereas the drive/excitation tone is swept in frequency. This technique, known simply as two-tone spectroscopy, is standard in the field [9], and is conceptually simple in the idealized case in which the drive is known to implement a single $\pi$ pulse and state readout using the monitor tone is done in a time-resolved fashion. The base temperature of the dilution refrigerator naturally prepares the qubit in its ground state, and if the drive tone is resonant with the qubit transition, then the qubit will be excited, causing the homodyne signal at the monitor frequency to switch to a different amplitude and phase.
In particular, we use a fully continuous-wave protocol in which both the drive and monitor tones are applied continuously. In this case, if the drive tone is resonant with the qubit and strong compared to the qubit dissipation rate, $\gamma$, the qubit will be driven to the maximally mixed state. The measured microwave signal at the monitor frequency is thus an average of the values that would be obtained for a qubit purely in the |$\ket{g}$ state and that of a qubit in the |$\ket{e}$ state. The maximum achievable signal with continuous-wave two-tone spectroscopy is thus half of what is achievable in pulsed operation. However, this slight loss in signal can be vastly outweighed by the much greater robustness of the technique. For example, pulsed two-tone spectroscopy is very sensitive to the strength of the applied drive, and if the drive is too strong, it can apply an accidental $n\times 2\pi$ pulse, and the qubit transition will no longer yield any signal. In continuous-wave all reasonably strong drives will saturate the qubit to the maximally mixed state. This drive-power insensitivity is particularly advantageous for looking at the photonic bound states. Because these are hybrid polariton states whose qubit-like character changes rapidly with detuning near the band edges, the effective Rabi rate induced by the drive can likewise change rapidly with frequency.
@@ III.2 Coplanar Waveguide Lattices
Since superconducting microwave resonators can be defined lithographically with low disorder and high quality factors [11, 13], it is relatively straightforward to produce large resonator arrays with $50$ or more sites [10, 13, 14, 18], and there are many different resonator geometries to choose from when designing an array, each with different advantages or disadvantages. Lumped-element resonators, for example, are extremely compact [13, 14], which is favorable for large-scale arrays.
@@
In this work, we concentrate solely on the case of CPW resonator arrays because they have a series of properties which are ideal for creating novel waveguide QED systems: first, they are known to be fabricable at large-scale with low disorder [11, 10]; second, their one-dimensional nature and compatibility with three-way couplers leads to photonic band structures which host linear band crossings, Dirac cones, and flat bands [18, 12]; and third, the frequency and effective hopping rate of a CPW resonator is insensitive to the shape of the resonator. This last property makes it possible for CPW arrays to realize resonator packings which do not obey the conventional rules of two-dimensional crystallography, allowing them to realize lattices in effective curved spaces [18] or giving the flexibility to fit a chosen lattice into a standard chip package. For example, a quasi-one-dimensional lattice with translation symmetry only in one direction, such as the one we study here (and discuss in detail in Sec. IV), would require a very long rectangular chip if made using CPW resonators of uniform size and shape. However, as shown in Fig. 2, it can be fabricated on a square chip by using a switchback pattern of unit cells, and the individual unit cells can be made compact by making use of resonators of different shapes.
The circuit theory of CPW lattices and the effective tight-binding description for photons in these lattices have been worked out in detail in Refs. [58, 59, 46, 12], so we will not provide a full treatment and set of derivations here. We instead summarize the key aspects of the CPW resonator modes and coupling structure necessary to show that resonator lattices can give rise to two distinct band structures and to see the origins of beyond tight-binding corrections, which begin to be relevant for the hopping strengths present in our device. The key results outlined below closely follow the treatments of Refs. [46, 12].
III.2.1 Single CPW Resonators

### v2

@@ III.1 Transmon Qubits and Readout
[…]
Because of the low transition frequencies of transmon qubits, they cannot be read out via fluorescence or photon absorption, as would be standard for an atomic qubit. Instead, standard practice is to exploit off-resonant coupling to a microwave resonator with frequency $\omega_{r}$ to perform readout [Blais:revmodphys]. In the dispersive regime, where $|\omega_{q}-\omega_{r}|\gg g$, and approximating the transmon as a two-level system, the effective Hamiltonian takes the form:
@@
where $\Delta=\omega_{q}-\omega_{r}$ and $\omega_{q}^{\prime}=\omega_{q}+g^{2}/\Delta$ is the Lamb-shifted qubit frequency. Crucially, the term $g^{2}\sigma_{z}a^{\dagger}a/\Delta$ constitutes a qubit-state dependent frequency shift of the resonator, known as the dispersive shift. For a more detailed description of corrections to the magnitude of the dispersive shift due to higher energy levels of the transmon, see Ref. [Blais:revmodphys]. Due to the native separation of energy scales between the coherent coupling strength $g$ of a transmon and the native decay rates, transmons can easily reach the so-called strong dispersive regime in which the dispersive shift is comparable to or exceeds the resonator linewidth $\kappa$. In this regime, the amplitude and phase of a monitor tone impinging on the resonator are strongly affected by the state of the qubit, and, as a result, homodyne or heterodyne measurements of this microwave tone yield QND measurements of the qubit state [Blais:revmodphys].
@@
Many geometries of resonator have been used in conjunction with transmons; however, in this work we concentrate only on the case of capacitive coupling to CPW resonators [Goppl_2008], because these resonators have the demonstrated capability to implement a large variety of unconventional band structures [Kollar:2019hyperbolic, Kollar:2019linegraph]. These systems routinely display $g/2\pi$’s in the range of 5050300300 MHz for single CPW resonators. In the case of an array of CPW resonators, such as the photonic crystal studied in this paper, the natural photonic degrees of freedom are normal modes, rather than individual cavity modes. In an array of $N$ resonators, a typical normal mode will be delocalized over the entire device, and due to the $N$-fold increase in the mode volume over that of a single isolated resonator, the light-matter coupling will become $g_{\mathrm{norm}}\sim g_{0}/\sqrt{N}$, where $g_{0}$ is the coupling between the qubit and the particular resonator in which it is located. Therefore, a CPW array with $\mathcal{O}(50-100)$ resonators can still have dispersive shifts which are an appreciable fraction of a MHz at a detuning of a GHz. Hence, the dispersive readout techniques which are normally used on transmon qubits in single-mode resonators can also be applied in resonator-array photonic crystal systems, using the normal modes.
@@ III.2 Coplanar Waveguide Lattices
Since superconducting microwave resonators can be defined lithographically with low disorder and high quality factors [Houck:earlylattice, Painter_Markov], it is relatively straightforward to produce large resonator arrays with $50$ or more sites [Underwood:imaging, Painter_Markov, Painter_SSH, Kollar:2019hyperbolic, Fitzpatrick_2017], and there are many different resonator geometries to choose from when designing an array, each with different advantages or disadvantages. Lumped-element resonators, for example, are extremely compact [Painter_Markov, Painter_SSH], which is favorable for large-scale arrays.
@@
In this work, we concentrate solely on the case of CPW resonator arrays because they have a series of properties which are ideal for creating novel waveguide QED systems: first, they are known to be fabricable at large-scale with low disorder [Houck:earlylattice, Underwood:imaging]; second, the frequency and effective hopping rate of a CPW resonator is insensitive to the shape of the resonator [Kollar:2019hyperbolic]; and third, their one-dimensional nature and compatibility with three-way couplers leads to photonic band structures which host quadratic bands, Dirac cones, and flat bands [Kollar:2019hyperbolic, Kollar:2019linegraph].
Flat bands, which are highly unusual in naturally occurring band structures, are ubiquitous in CPW lattices because the individual lattice sites are one-dimensional line-like objects which connect at their end points, rather than point-like objects. As a result, the lattices produced correspond to a special class of graphs known as line graphs [Kollar:2019linegraph, Kollar:2019hyperbolic], which are generalizations of the kagome lattice. Every line graph originate from an underlying root graph, for example, the kagome lattice is the line graph of graphene, and all the properties of the line graph can be derived from the underlying root graph [Kollar:2019linegraph]. The mapping from the root band structure to the line-graph band structure for both symmetric and asymmetric on-site wavefunctions was worked out in Ref. [Kollar:2019linegraph], and both cases are guaranteed to have flat bands.
The insensitivity of CPW resonator frequencies and hopping rates to resonator shape allows the design of CPW lattices to be extremely flexible, making it possible for CPW arrays to realize resonator packings which do not obey the conventional rules of two-dimensional crystallography, making it possible to realize lattices in effective curved spaces [Kollar:2019hyperbolic] or giving the flexibility to fit a chosen lattice into a standard chip package. For example, a quasi-one-dimensional lattice with translation symmetry only in one direction, such as the one we study here (and discuss in detail in Sec. IV), would require a very long rectangular chip if made using CPW resonators of uniform size and shape. However, as shown in Fig. 2, it can be fabricated on a square chip by using a switchback pattern of unit cells, and the individual unit cells can be made compact by making use of resonators of different shapes.
In the remainder of this section, we briefly summarize the theory underlying CPW resonator lattices. A more detailed description can be found in Appendix A. An isolated CPW resonator of length $L$ bounded by capacitors has cosinusoidal standing-wave eigenmodes with antinodes at both ends of the resonator. The corresponding mode functions are
$\bar{\Phi}_{\mu}(x)\propto\cos\left(\frac{\mu\pi x}{L}\right)\!,$ (8)
with one end of the resonator taken to be at $x=0$ for simplicity. The integer $\mu$ labels the different harmonics of a single resonator, with the lowest-frequency, fundamental mode corresponding to $\mu=1$. This antisymmetric fundamental mode, with opposite values of $\bar{\Phi}$ at each end of the resonator, satisfies $L=\lambda/2$ and is thus known as the half-wave mode. The second-harmonic mode, corresponding to $\mu=2$, is symmetric and occurs at twice the frequency of the half-wave mode. It satisfies $L=\lambda$, and is known as the full-wave mode. A capacitively coupled transmon at the end of the resonator couples to an antinode of all modes $\mu$, and the strength of the qubit-mode coupling, $g_{\mu}$, scales as $\sqrt{\omega_{\mu}\omega_{q}}$, increasing with higher mode numbers (see Appendix A.4 for details).
The resonators in the CPW lattice presented in this work are butt-coupled through 3-way coupling capacitors, previously developed in Refs. [Houck_Nature_QS, Houck:earlylattice, Underwood:imaging, Kollar:2019hyperbolic], where the center pin of three resonators end in close proximity to one another. The capacitance between neighboring resonators, $C_{c}$, introduces a coupling term to the Hamiltonian, and the resultant CPW lattice can be well described as an effective photonic tight-binding model of the form [Koch:AnnPhysBerl, Koch:TRS_breaking]
$H_{\mu}=\sum_{n}\hbar\omega_{\mu}a_{n}^{{\dagger}}a_{n}-\sum_{\langle n,n^{\prime}\rangle}\hbar t^{(\mu)}_{n,n^{\prime}}a_{n}^{{\dagger}}a_{n^{\prime}}.$ (9)
Here, $\omega_{\mu}$ is the resonant frequency of a single resonator, $\langle n,n^{\prime}\rangle$ indicates neighboring resonators that share a coupler, and $t^{(\mu)}_{n,n^{\prime}}$ is the effective hopping between those resonators.
The sign of the hopping in CPW lattices is generally unusual, with long-wavelength normal modes occurring at high energy and short-wavelength modes occurring at low energy [Houck:earlylattice, Koch:AnnPhysBerl, Kollar:2019linegraph], for example, the kagome-like flat-band induced by the line-graph structure is always at the bottom of the spectrum [Kollar:2019linegraph]. One way to understand this is that in the standard black-box quantization circuit picture, the coupling capacitor is equivalent to a modulation in the kinetic energy or the effective mass, rather than a potential energy barrier (see Ref. [Koch:TRS_breaking]).